the number of homes built can be modeled by an exponential function,
H= p * a(t) , where p is the number of new homes built in the first year recorded.
p=10 t =3 a=2
im not sure how to caculate it.

Answer by rapaljer(4551)

H=p*a^t
H=10*2^3
H=10*8=80 homes

R^2